OpenStudy (anonymous):

If 1.0 gram of matter was converted to energy, how much energy would be formed? A. 9.0 x 10^-13 joules B. 9.0 x 10^13 joules C. 3.0 x 10^13 joules D. 3.0 x 10^8 joules Chemistry help??

OpenStudy (abhisar):

\(\sf Energy = mass \times (speed~of~light)^2\) Energy = \(\sf 10^{-3}Kg \times (3 \times 10^8m/s)^2=? \)